Solve by the elimination method.
3x-9y=3
2x-9y=4
WHat is the solution to the system?
A) The solution is ___( type an ordered pair. Use integers or fractions for any numbers in the expression)
B) There are infinitely many solutions.
c) There is no solution

Answer by brucewill(101)

By "elimination", we mean eliminate one variable, and then solve for the other.
In the case, we can simply subtract the second statement from the first producing: x = -1.
Now, we can substitute the -1 into either equation. I'll choose the first:
-3 -9y = 3
-9y = 6
y = -(2/3).
The answer then is (-1, -2/3)
You can substitute for x & y back into the original equations:
3(-1) - 9(-2/3) = 3, -3 + 6 = 3, 3 = 3. OK
2(-1) - 9(-2/3) = 4, -2 + 6 = 4; 4 = 4. OK.